# Prospects for High-Frequency Gravitational-Wave Detection with GEO600

### v1

@@ A.1 Unfolded Michelson Interferometer
where the unit vectors ^$\hat{\ell},\hat{m},\hat{n}$ form a right-handed orthonormal basis. In this discussion, we suppress the relative rotation that is due to the polarization angle. In the long-wavelength approximation, the Michelson interferometer response is given by
@@
$D(\hat{a},f)=\frac{\mathrm{e}^{-\mathrm{i}2\pi fT}}{2}\!\left[\mathrm{e}^{\mathrm{i}\pi fT_{+}}\operatorname{sinc}\!\left(\pi fT_{-}\right)+\mathrm{e}^{-\mathrm{i}\pi fT_{-}}\operatorname{sinc}\!\left(\pi fT_{+}\right)\right],$ (25)
and define $T_{\pm}=T(1\pm\hat{a}\cdot\hat{n})$. In the end, the frequency-domain GW signal is
@@ A.2 Folded Michelson Interferometer
Figure 7: Left: sky-averaged antenna patterns $\mathcal{R}(f)$ for aLIGO (light blue), folded GEO600 (purple), a 600m Michelson interferometer (brown), and a 1200m Michelson interferometer (gray). The folding factor given in Eq. (31) relates the angular response of an unfolded Michelson interferometer (brown) to its folded counterpart (purple); one can see that the folding factor sharply reduces the angular response at frequencies where it was previously strong. Right: antenna patterns $F(\theta_{\hat{n}},\phi_{\hat{n}})$ of folded GEO600 for the following incident-angled GWs: overhead (light blue), completely along one detector arm (gray), and two in-between angles (purple and brown). Here, $F(\theta_{\hat{n}},\phi_{\hat{n}})=\sqrt{|F_{+}(f;\theta_{\hat{n}},\phi_{\hat{n}})|^{2}+|F_{\times}(f;\theta_{\hat{n}},\phi_{\hat{n}})|^{2}}$ since we assume an inclination angle of zero. One can see that the angular behavior of antenna patterns is very frequency dependent at high frequencies.
@@ A.3 Sky Averaging
where $F_{+}=F_{+}(\theta_{\hat{n}},\phi_{\hat{n}})$ only at $\psi=0$. If we average over the direction that the GW arrives, we can write this as
@@
[…]
@@ B.4 Laser Frequency Noise
Laser frequency noise is created by frequency fluctuations of the main laser light while traveling through the interferometer. The ASD for the raw frequency noise is given in dimensions of Hz/$\text{Hz}/\!\sqrt{\text{Hz}}$ by [140] as
@@ B.6.2 Test Mass Substrate Thermal Noise
$A_{x}^{\text{TMTE}}(\omega)=\sqrt{\begin{aligned} &\left[1+\frac{k_{s}r_{0}}{\sqrt{2\pi}}\right]\times\\ &\left[\frac{8\kappa\alpha^{2}\left(1+\sigma\right)^{2}k_{B}T_{0}^{2}}{\sqrt{2\pi}\rho^{2}C_{p}^{2}r_{0}^{3}\omega^{2}}\right]\end{aligned}}.$ (55)
@@ B.6.3 Test Mass Coating Thermal Noise
$\mathbb{E}_{\text{Coating}}=\frac{HF_{0}^{2}}{4\pi{r_{0}^{2}}}\biggl{[}\frac{(1+\sigma_{c})(1-2\sigma_{c})}{Y_{c}(1-\sigma_{c})}$ (57)
@@ B.6.4 Beam Splitter Thermal Noise
$A_{x}^{\text{BSTR}}(\omega)=\!\!\!\sqrt{\begin{aligned} &\!\!\left[\frac{4{k_{B}}{\kappa}{T^{2}}{\beta^{2}}a^{\prime}\left(\eta+\eta^{-1}\right)}{\pi\left(C{\rho}{r_{0}}^{2}\omega\right)^{2}2\eta^{2}}\right]\times\\ &\!\!\left[1+\frac{2{k^{2}}{r_{0}}^{2}\eta}{\left(\eta+\eta^{-1}\right)\left(1\!+\!\left(2kl_{th}(\omega)\right)^{4}\right)}\right].\end{aligned}}$ (59)

### v2

@@ A.1 Unfolded Michelson Interferometer
where the unit vectors $\hat{\ell},\hat{m},\hat{n}$ form a right-handed orthonormal basis. In this discussion, we suppress the relative rotation that is due to the polarization angle. In the long-wavelength approximation, the Michelson interferometer response is given by
@@
$\hskip-10.00002ptD(\hat{a},f)=\frac{\mathrm{e}^{-\mathrm{i}2\pi fT}}{2}\!\left[\mathrm{e}^{\mathrm{i}\pi fT_{+}}\operatorname{sinc}\!\left(\pi fT_{-}\right)+\mathrm{e}^{-\mathrm{i}\pi fT_{-}}\operatorname{sinc}\!\left(\pi fT_{+}\right)\right],$ (25)
[…]
@@ A.2 Folded Michelson Interferometer
[…]
@@ A.3 Sky Averaging
[…]
@@
$=\frac{1}{4\pi}\int d\Omega_{L}\left(A_{+}^{2}(\iota)+A_{\times}^{2}(\iota)\right),$ (38)
@@ B.4 Laser Frequency Noise
Laser frequency noise is created by frequency fluctuations of the main laser light while traveling through the interferometer. The ASD for the raw frequency noise is given in dimensions of $\text{Hz}/\!\sqrt{\text{Hz}}$ by [141] as
@@ B.6.2 Test Mass Substrate Thermal Noise
$\hskip-5.0ptA_{x}^{\text{TMTE}}(\omega)=\sqrt{\begin{aligned} &\left[1+\frac{k_{s}r_{0}}{\sqrt{2\pi}}\right]\times\\ &\left[\frac{8\kappa\alpha^{2}\left(1+\sigma\right)^{2}k_{B}T_{0}^{2}}{\sqrt{2\pi}\rho^{2}C_{p}^{2}r_{0}^{3}\omega^{2}}\right]\end{aligned}}.$ (55)
@@ B.6.3 Test Mass Coating Thermal Noise
$\hskip-5.0pt\mathbb{E}_{\text{Coating}}=\frac{HF_{0}^{2}}{4\pi{r_{0}^{2}}}\biggl[\frac{(1+\sigma_{c})(1-2\sigma_{c})}{Y_{c}(1-\sigma_{c})}$ (57)
@@ B.6.4 Beam Splitter Thermal Noise
$\hskip-5.0ptA_{x}^{\text{BSTR}}(\omega)=\!\!\!\sqrt{\begin{aligned} &\!\!\left[\frac{4{k_{B}}{\kappa}{T^{2}}{\beta^{2}}a^{\prime}\left(\eta+\eta^{-1}\right)}{\pi\left(C{\rho}{r_{0}}^{2}\omega\right)^{2}2\eta^{2}}\right]\times\\ &\!\!\left[1+\frac{2{k^{2}}{r_{0}}^{2}\eta}{\left(\eta+\eta^{-1}\right)\left(1\!+\!\left(2kl_{th}(\omega)\right)^{4}\right)}\right].\end{aligned}}$ (59)